OpenStudy (barrelracer011):

When 27 times x squared times z all over negative 3 times x squared times z to the fourth power is completely simplified, the exponent on the variable z is _____.

OpenStudy (barrelracer011):



OpenStudy (barrelracer011):

@johnweldon1993

OpenStudy (barrelracer011):

I think its -3

OpenStudy (johnweldon1993):

And you would be correct since when you divide exponents with the same base, you just subtract the exponents...so yes \[\large \frac{z^1}{z^4} = z^{1 - 4} = z^{-3}\]

OpenStudy (barrelracer011):

Thanks! I have one more can you help?

OpenStudy (johnweldon1993):

Sure!

OpenStudy (barrelracer011):

Choose the correct simplification of the expression 9 over h to the power of negative 3. h to the 3rd power over 9 9 over h to the 3rd power 9h3 This expression is already simplified.

OpenStudy (barrelracer011):



OpenStudy (johnweldon1993):

Any ideas first?